# In Search of the Long-Tail: Systematic Generation of Long-Tail Inferential Knowledge via Logical Rule Guided Search

### v2

@@ Abstract
State-of-the-art LLMs outperform humans on reasoning tasks such as Natural Language Inference. Recent works evaluating LLMs note a marked performance drop on input data from the low-probability distribution, i.e., the long-tail. Therefore, we focus on systematically generating statements involving long-tail inferential knowledge for more effective evaluation of LLMs in the reasoning space. We first propose a novel framework Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]) [^1] https://github.com/INK-USC/LINK/ that generates factually correct and long-tail knowledge statements grounded on symbolic rule templates; LINK effectively generates data in the long-tail distribution that zero-shot prompted LLMs are unable to reach, and outperforms zero-shot GPT4 on factual correctness by 5%. We further use the data generated by LINK to construct a dataset Logic-Induced-Long-Tail (LINT) [^2] https://doi.org/10.5281/zenodo.10691770 that can be used to evaluate downstream models on the long-tail distribution; LINT contains 108K knowledge statements spanning four domains. We use LINT to test LLMs on an entailment classification task and find that model performances drop by as high as 5% in the long-tail distribution compared to head distribution. Our work shows the utility of evaluating models in the long-tail distribution, and calls for more research on generating evaluation data in the long-tail distribution.
@@
[…]
Figure 1: An illustration of the distribution of statements generated by different methods, after being filtered on factual correctness by human evaluation. For the same symbolic rule, all methods generate the same number of statements, but LINK(Refer to caption) has more correct statements than zero-shot ChatGPT and zero-shot GPT4(Refer to caption).  LINK long-tail generations clearly lie in a more long-tail part of the probability distribution than the zero-shot GPT4 generations.
@@ 1 Introduction
Large language models (LLMs) have approached broadly human-level performance on an increasing number of benchmarks (OpenAI, 2023; Ouyang et al., 2022). The problem of evaluating models that potentially outperform (non-expert) humans on most skills relevant to the task seems more imminent than ever (Bowman et al., 2022). Natural Language Inference (Bowman et al., 2015; Zellers et al., 2019), which requires extensive knowledge and complex reasoning about entities and events, is one of such tasks on which LLMs have impressive performance (Achiam et al., 2023; Touvron et al., 2023b; Jiang et al., 2023). Recent works on evaluating LLMs on challenge reasoning sets notice a marked decline in performance when facing input data from the “long-tail”, i.e., the low-probability part of the natural language’s distribution [^3] See Section 2.1 for discussion on the relationship between the probability distribution of natural language and LLMs.(McCoy et al., 2023; Razeghi et al., 2022).
To have a well-rounded evaluation of LLMs within the reasoning space, one needs access to inferential knowledge statements (Figure 1) collected systematically from the long-tail to form challenging evaluation sets (i.e., long-tail inferential knowledge). However, so far there has not been any systematic framework that targets generating data in the long-tail of probability distribution of language models. Crowdsourcing long-tail data is difficult because of human cognitive bias (Tversky and Kahneman, 1973, 1974), and LLMs’ generation in the long-tail distribution is hindered by their pretraining task of “most likely" next token (McCoy et al., 2023).
In this paper, we propose a logical rule-guided long-tail knowledge generation framework, Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]), for generating inferential knowledge statements that are both factually correct and in the long-tail  – a task that is fundamentally challenging to SoTA LLMs. To achieve the two goals simultaneously, we first ground the generation of inferential knowledge statements with symbolic rules, which consist of variables and predicates that specify the relationship between variables. The knowledge statements are ensured to be valid as long as the filled values of the variables satisfy all predicates in the rule. Doing so significantly simplifies the task from sentence generation to value search with the given rules. Specifically, we propose knowledge beam search (Figure 2) to utilize the modularity of symbolic rules by searching for values of the variables one by one in a designated order – an algorithm inspired by sequential beam search. At each step, we prompt an LLM for plausible values that satisfy all predicates in which the variable appears. The step-wise search design also allows for granular control of the statement’s correctness and likelihood as it incorporates a critic model that checks for data type conformity and factual correctness and a reranker model that ensures that the values fall into the long-tail of the distribution.
LINK shows clear advantage in generating long-tail knowledge statements compared to prompting LLMs (Figure 1). We find that even with symbolic rules presented in the prompt, e.g., by instructing LLM to fill in values for the rules, the prompt-based methods using ChatGPT (gpt-3.5-turbo) and GPT4 (gpt-4) still struggle to generate knowledge statements that are both correct and in the long-tail distribution. This implies that the prompt-based methods may suffer from directly retrieving long-tail knowledge from the natural language distribution. In addition, human annotators find that compared to LINK, the statements generated by prompt-based methods are prone to have more factual errors. Further analysis shows (1) the key contribution to the long-tail generation quality comes from the LINK pipeline, (2) LINK can be effectively and flexibly applied across different knowledge models, and (3) post-hoc reranking does not reach into the long-tail distribution as LINK.
Using LINK, we curate a dataset, Logic-Induced-Long-Tail (LINT), which contains 108k knowledge statements spanning across four different domains, presented in a premise, conclusion format (Table 1). LINT can be a useful resource for the community to more thoroughly test LLMs’ capabilities in the long-tail distribution. We use a subset of LINT to create an evaluation task in the form of entailment classification, where LLMs are asked to judge true or false on a knowledge statement. In the long-tail test set, GPT4’s capability in identifying incorrect knowledge drop by as high as 3% with zero-shot prompting, and as high as 4% with Chain-of-Thought prompting (Wei et al., 2022). GPT4’s performance gap from human baseline on the long-tail set is as high as 19%. In addition, we probe models with different template formats and show that template bias decreases in the long-tail distribution.
@@
Table 1: Examples of inferential knowledge in each domain of LINT, in premise,conclusion format.
Locational
Premise: Person X is located at Long Beach. Conclusion: Person X can plant Palms.
Capability and Advice
Premise: Person X has Asthma. Conclusion: Person X should take Terbutaline.
Temporal
Premise: Plant X vanished in Safavid. Conclusion: Plant X cannot be used to conceal MAS 36.
Natural Properties
Premise: Drawer X has trouble containing Scissors. Conclusion: Router cannot be placed in Drawer X.
LINK is our first step towards the generation of long-tail data. The biggest challenge of generating long-tail data for LLMs is juggling enforcing factual correctness with remaining in lower data likelihood, and our logic-induced knowledge beam search framework provides a satisfying solution to this problem. Our work serves as a starting point of the series of research on systematically generating data in the long-tail distribution, and our dataset is a useful resource for evaluating LLMs in the long-tail distribution.
@@
Figure 2: Overview of knowledge beam search (§ 2.3), consisting of a knowledge model, a critic model and a reranker model. We populate the symbolic rule with values and convert it into a natural language statement. This example demonstrates searching B conditioned on the values of A and Z from previous steps. Note that we only verbalize the predicates containing Person X, because all other predicates contain knowledge that the model should have.
@@ 2 Logic-Induced-Knowledge-Search (LINK[Uncaptioned image])
2.1 Long-tail Knowledge Statements
Admittedly, no single language model can resemble the true natural language distribution. However, in Appendix D.2, we show that the long-tail distribution of different language models overlap, and that this evidence supports our assumption that a universal natural language distribution exists, and that one language model can approximate the long-tail distribution of other language models.
Therefore, we approximate the written natural language distribution using the existing largest model that outputs log probability, InstructGPT (text-davinci-003) [^4] Since InstructGPT was deprecated on 2024/01/04, we used llama2-70b for distribution approximation in later studies. See Appendix D.2 for their comparison.. Although it is hard to define the exact boundary between the long-tail and the head distribution, we can define them comparatively: with similar length and format, statements in the long-tail distribution have lower likelihood than those in the head distribution. When we generate long-tail statements, we aim to generate those landing in the lower likelihood distribution, compared to other statements of similar length and format.
Inferential knowledge statements are commonly represented under a premise, conclusion structure (Sap et al., 2019). The premise introduces factual knowledge about a person or an object, and the conclusion is some information about the person or the object deducible from the premise.
Intuitively, for statements with similar length and format, the likelihood of content words in the statements heavily impacts the likelihood of the statements. Therefore, we can create long-tail knowledge statements from a symbolic rule and then fill in content words that are less likely to appear in the training corpus. Creating knowledge statements from symbolic rules has two benefits: (1) Symbolic rules are designed to be correct, so we alleviate the pressure of ensuring the deductive plausibility of the statement throughout the entire generation process. (2) The generation process can be broken down into multiple steps, each of which is conditioned on only one variable. The knowledge generation task will be much easier for the model, and it is easier to maneuver the distribution of individual values than the entire sentence.
@@ 2.2 Curating Symbolic Rules
A symbolic rule consists of a premise and a conclusion. The conclusion is a single predicate, while the premise contains a set of predicates connected by &\&& operators. Each predicate is a triple of a verb phrase, a subject and an object, and each variable in the symbolic rule has a designated data type. See example below:
allergic_to(Person X, Allergen A) & one_type_of(Ingredient Z, Allergen A)
& ingredient_in(Ingredient Z, Dish B) $\to$ cannot_eat(Person X, Dish B)
A symbolic rule must follow some constraints to turn into a knowledge statement:
The premise and conclusion need to have the same subject.
Different subjects in the premise and conclusion make the rule more fuzzy and less verifiable.
The symbolic rule should be linearly chained.
Starting from the subject in the premise, we should be able to find a chain of variables that connects the subject to the object in the conclusion (or vice versa), and each variable should appear once and only once. This is to simplify the search process.
The premise and conclusion should not be paraphrases.
@@
We curated symbolic rules that span across four domains: temporal, locational, capability and advice, and natural properties, totaling 149 person-related rules and 268 object-related rules. Details of our symbolic rules are in Appendix A.
@@ 2.3 Knowledge Beam Search
Since all variables are linearly chained, we can search them one by one without repetition. We always start with the subject of the sentence – the person or the object, represented as Datatype X. In the rule in § 2.2, for example, we start with Person X in the premise and find a chain of variables that connects it to the object in the conclusion: X, A, Z, B.
@@
For each variable, we construct a prompt using all predicates that contain that variable and other previously searched variables. For example, to search variable B in the rule in § 2.2, we include predicate ingredient_in(Ingredient Z, Dish B). We assume Z=butter and construct the prompt as follows:
@@
Prompting InstructGPT for knowledge.
For each partially searched beam, we obtain 200 values of the current variable from the knowledge model. We call OpenAI API 4 times, generating 50 values each time (temperature=0.7  [^5] We keep top_p=1 for maximum diversity, and top_k is hidden in the OpenAI API. Ablation on temperature is in Appendix C.1.). After each call, we verify the values using a critic model (see paragraph below). To prevent duplicates, we explicitly instruct the model not to generate verified correct values and modify the token probability of incorrect values. We implement an early stop mechanism: if for two consecutive calls we do not get any correct values, we terminate the search for the beam.
@@
We use huggingface default implementation of Flan-T5-XXL (Chung et al., 2022), an instruction-tuned model that can be used zero-shot, as the critic that checks data type conformity and factual correctness of the values. We ask the model to output yes/no on the correctness of a given statement. For data type conformity, the statement is “{value} is a {data type}.” For factual correctness, we convert the symbolic predicate into a natural language statement. We obtain the yes token probability and dynamically adjust the threshold for accepting values for different predicates. For details of the critic model, see Appendix B.
Pushing values to long-tail distribution.
At each search step, we convert symbolic predicates into natural language statements (ingredient_in(butter, saag chicken) normal-$\to$ Butter is an ingredient in saag chicken) and concatenate them with “and”. We obtain the sentence likelihood using huggingface default implementation of llama-7B (Touvron et al., 2023a)666llama2 was not released at the time of experiments. and rank the sentences from the lowest likelihood to the highest likelihood. We take top the 75% values unless there are more than 200 values, in which case we take the top 200 values. Then we move on to the next variable. When searching for values in the head distribution, we rank the sentences from the highest to the lowest likelihood.
From 149 person-related rules and 268 object-related rules across four domains, we curate our dataset Logic-Induced-Long-Tail(LINT) that consists of 54K long-tail knowledge statements. In addition, we also release 54K head distribution statements that are also searched with the LINK framework. Domain-wise statistics of LINT can be found in Appendix F.
3 Can LLM generate in the Long-tail Distribution?
3.1 Generating Long-tail Statements from Symbolic Rules
Prompt
In the following sentence, A is a Food allergen, B is a Name of a dish or food, Z is a Ingredient. Find values of A, B, Z to fill in the blank in the sentence ‘If Person X is allergic to [A] and [ Z] is a ingredient in [ B] and [Z] is one type of [ A], then Person X is not able to eat [ B].’ and make it a grammatical and correct sentence. Give me 50 values in the format ‘1. A=, B=, Z=’.

### v3

@@ Abstract
To effectively use large language models (LLMs) for real-world queries, it is imperative that they generalize to the long-tail distribution, i.e., rare examples where models exhibit low confidence. In this work, we take the first step towards evaluating LLMs in the long-tail distribution of inferential knowledge. We exemplify long-tail evaluation on the Natural Language Inference task. First, we introduce Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]), a systematic long-tail data generation framework, to obtain factually-correct yet long-tail inferential statements. LINK uses variable-wise prompting grounded on symbolic rules to seek low-confidence statements while ensuring factual correctness. We then use LINK to curate Logic-Induced-Long-Tail (LINT), a large-scale long-tail inferential knowledge dataset that contains 108K statements spanning four domains. We evaluate popular LLMs on LINT; we find that state-of-the-art LLMs show significant performance drop (21% relative drop for GPT4) on long-tail data as compared to on head distribution data, and smaller models show even more generalization weakness. These results further underscore the necessity of long-tail evaluation in developing generalizable LLMs. [^1] https://github.com/INK-USC/LINK
@@
Huihan Li1     Yuting Ning2     Zeyi Liao2     Siyuan Wang3     Xiang Lorraine Li4,5 Ximing Lu4,6     Wenting Zhao7     Faeze Brahman4     Yejin Choi4,6     Xiang Ren1,4 1University of Southern California   2 Ohio State University   3Fudan University 4Allen Institute for Artificial Intelligence   5 University of Pittsburgh 6 University of Washington   7 Cornell University {huihanl,xiangren}@usc.edu
Figure 1: Overview of our motivation, long-tail data generation framework, and model evaluation.
@@ 1 Introduction
Generalization, especially to unfamiliar and novel situations, is a cornerstone for the usability of large language models (LLMs) in addressing varied real-world inquiries. This imminent demand necessitates evaluation of LLMs in the long-tail  distribution (the space consisting of unfamiliar examples on which the model has low confidence). Previous works, mostly focusing on model memorization issue, define long-tail knowledge using the frequency of entities in a knowledge base (Cao et al., 2020), in the pre-training dataset (Kandpal et al., 2023), or in Wikipedia search (Mallen et al., 2022). Godbole and Jia (2022) introduces a general definition for long-tail statements, where long-tail examples are assigned lower likelihood by a pre-trained language model. We follow this definition which applies to various data format and language task – for any set of statements with similar length and format, those in the long-tail distribution cannot be generated or are generated with low confidence by the models, compared to those in the head distribution.
Recent works have noticed that LLMs have a marked decline in performance when facing inputs from the long-tail  McCoy et al. (2023); Razeghi et al. (2022). Hallucination, for example, is found to be correlated with data being in the long-tail distribution (Li et al., 2024; Yu et al., 2024). LLMs’ ineffective utilization of long-tail knowledge impacts its reasoning capabilities and raises reliability concerns in downstream tasks (Huang et al., 2023).
Evaluation in the long-tail distribution requires systematic generation of long-tail data. However, obtaining examples in the long-tail is non-trivial. With state-of-the-art LLMs being trained on vast volume of data on the internet (OpenAI, 2023; Touvron et al., 2023b), it is increasingly difficult to find unseen examples that can effectively test model generalization to its low-confidence end. Crowdsourcing long-tail data is also difficult because of human cognitive bias (Tversky and Kahneman, 1973, 1974), and LLMs’ generation in the long-tail distribution is hindered by their pretraining task of “most likely" next token (McCoy et al., 2023).
While demonstrating long-tail evaluation across all applications and domains is not feasible within the scope of one paper, this work focuses on inferential knowledge statement in the form of Natural Language Inference (NLI) task (Bowman et al., 2015; Zellers et al., 2019): NLI requires extensive knowledge and complex reasoning about entities and events, and is one of such tasks on which LLMs have impressive performance (Achiam et al., 2023; Touvron et al., 2023b; Jiang et al., 2023). Following Sap et al. (2019), we structure inferential knowledge as if-then relations with variables, written in a premise, conclusion format (Table 1).
First, we make long-tail inferential knowledge generation possible. We propose a novel and lightweight long-tail inferential knowledge generation framework, Logic-Induced-Knowledge-Search (LINK[Uncaptioned image]) (§ 2), a variable-wise prompting framework grounded on symbolic rules. This framework enables us to obtain long-tail knowledge statements from existing LLMs. Our evaluation shows that by taking simply instructions ChatGPT(gpt-3.5-turbo) and GPT4(gpt-4) can only produce statements in the head distribution that also have lower factual correctness, but using LINK with the LLMs improves on both distribution conformity and factual correctness (§ 3).
@@
Table 1: Examples of inferential knowledge in each domain of LINT, in a premise (P), conclusion (C) format.
Locational Head P: Organization X has a branch in Great Lakes Region. C: Organization X has office in North America.
Long-tail P: Organization X has a branch in Tarapaca Region. C: Organization X has office in South America.
Outcome and Effect Head P: Person X has Asthma. C: Person X should take Inhaled antiinflammatory drugs.
Long-tail P: Person X has Hepatitis. C: Person X should take Sofosbuvir.
Temporal Head P: Plant X vanished in Paleolithic Era. C: Plant X cannot surround Notre Dame de Paris.
Long-tail P: Plant X vanished in Classical Greece. C: Plant X cannot surround Belém Tower.
Natural Properties Head P: Bag X has trouble containing Clarinet. C: Upright Piano cannot fit in Bag X.
Long-tail P: Bag X has trouble containing Pandeiro. C: Dhak cannot fit in Bag X.
Second, we test LLMs’ long-tail generalization capability on data generated by LINK(§ 4). We produce a large-scale dataset, Logic-Induced-Long-Tail (LINT), which contains 108k knowledge statements spanning across 4 different domains(Table 1). In the long-tail test set of LINT, GPT4’s capability in identifying incorrect knowledge drop by 21% relative to the test set in the head distribution, and the gap is even larger for other models we tested (ChatGPT, llama2-70b). At the same time, human performance significantly outperforms LLMs in both distributions, and stays consistent between head and long-tail test set.
Our work is the first to propose a systematic framework that generates data in the long-tail distribution. Using NLI as an example, we show that generating data in the long-tail distribution is an effective way for curating evaluation examples for LLM generalization. Our work serves as a starting point for the series of research on long-tail data discovery and generation, and motivates the community to incorporate long-tail evaluation into model building pipelines.
@@
Figure 2: Overview of knowledge beam search (§ 2.3). We demonstrate searching B conditioned on the values of A and Z from previous steps. We only verbalize the predicates containing Person X in the final statement as all other predicates contain knowledge that the model should have.
@@ 2 Logic-Induced-Knowledge-Search (LINK[Uncaptioned image])
In this section, we first explain the advantages of generating inferential knowledge grounded on symbolic rules, then illustrate our process of curating symbolic rules, and lastly explain knowledge beam search, our novel variable-wise search pipeline.
2.1 Advantages of Generating Inferential Knowledge from Symbolic Rules
Due to the fact that LLMs are pretrained with the task of generating the “most likely" next token, it is fundamentally challenging for them to directly generate long-tail data through prompting that are factually correct and have low likelihood. Using symbolic rules to guide the generation of knowledge statements have three benefits: (1) Symbolic rules are designed to be correct, so we alleviate the pressure of ensuring the deductive plausibility of the statement throughout the entire generation process. (2) The generation process can be broken down into multiple steps, each of which is conditioned on only one variable. Generating for one variable at a time will be much easier for the model, and it is easier to manipulate the distribution of individual values than the entire sentence. (3) From one symbolic rule, one can get abundant combinations of variable values as long as they satisfy each predicate in the rule, making the generation process scalable.
@@ 2.2 Curating Symbolic Rules
A symbolic rule consists of a premise and a conclusion. The conclusion is a single predicate, while the premise contains a set of predicates connected by &\&& operators. Each predicate is a triple of a verb phrase, a subject and an object, and each variable in the symbolic rule has a designated data type.
While there are infinitely many ways to construct symbols rules, we create ours using the principles of Compatibility and Mutual Exclusivity. Compatibility refers to a scenario in which one or more events enables another event to occur. We construct the premise and conclusion such that if all predicates in the premise are true, then the predicate in the conclusion can occur. Mutual Exclusivity refers to a scenario in which two or more events or conditions cannot occur simultaneously. We construct the premise and conclusion such that if all predicates in the premise are true, then the predicate in the conclusion will not occur. To construct such conditions, we add constraints such as time, location, or outcomes to variables in the symbolic rules to make their interaction possible/desirable or impossible/undesirable. In other words, the conclusion describes an event (certain interaction between the variables) and the premise depicts some combination of conditions.
For example, a symbolic rule constrained by compatibility looks like:
exists(Person X, Location  A) & lives(Plant Y, Climate  B)
& has_climate(Location  A, Climate B)
$\to$ can_plant(Person X, Plant Y)
And a symbolic rule constrained by mutual exclusivity looks like:
exists(Person X, Time Period  A) & exists(Plant Y, Time Period  B) & is_much_later_than(Time Period  A, Time Period B) $\to$ cannot_plant(Person X, Plant Y)
Here are some additional properties of our symbolic rules:
No tautologies.
@@
We create symbolic rules that span across four domains (of constraint type): temporal, locational, outcome and effect, and natural properties, totaling 149 person-related rules and 268 object-related rules. More about symbolic rules in Appendix B.
@@ 2.3 Knowledge Beam Search
Since all variables are linearly chained, we can search them one by one without repetition. We always start with the subject of the sentence – the person or the object, represented as Datatype X. In the rule in Table 2, for example, we start with Person X in the premise and find a chain of variables that connects it to the object in the conclusion: X, A, Z, B.
@@
For each variable, we construct a prompt using all predicates that contain that variable and other previously searched variables. For example, to search variable B in the rule in Table 2, we include predicate ingredient_in(Ingredient Z, Dish B). We assume Z=butter and construct the prompt as follows:
@@
Prompting for knowledge.
For each partially searched beam, we obtain 200 values of the current variable from the knowledge model  [^2] We use text-davinci-003 but one can use any model.. We call OpenAI API 4 times, generating 50 values each time (temperature=0.7  [^3] We keep top_p=1 for maximum diversity, and top_k is unchangable. Ablation on temperature in Appendix D.1.). After each call, we verify the values using a critic model (see paragraph below). To prevent duplicates, we explicitly instruct the model not to generate verified correct values and set logit_bias=-100 for incorrect values. We implement an early stop mechanism: if for two consecutive calls we do not get any correct values, we terminate the search for the beam.
@@
We use huggingface default implementation of Flan-T5-XXL (Chung et al., 2022), an instruction-tuned model that can be used zero-shot, as the critic that checks data type conformity and factual correctness of the values. We ask the model to output yes/no on the correctness of a given statement. For data type conformity, the statement is “{value} is a {data type}.” For factual correctness, we convert the symbolic predicate into a natural language statement. We obtain the yes token probability and dynamically adjust the threshold for accepting values for different predicates. More about the critic model in Appendix C.
Pushing values to long-tail distribution with reranking.
At each search step, we convert symbolic predicates into natural language statements (ingredient_in(butter, saag chicken) $\to$ Butter is an ingredient in saag chicken) and concatenate them with “and”. We obtain the sentence likelihood using huggingface default implementation of llama-7B (Touvron et al., 2023a)444llama2 was not released at the time of experiments. and rerank the sentences from the lowest likelihood to the highest likelihood. We take top the 75% values unless there are more than 200 values, in which case we take the top 200 values. Then we move on to the next variable. To control data distribution during evaluation, we also generate statements in the head distribution, by ranking the sentences from the highest to the lowest likelihood.
From 149 person-related rules and 268 object-related rules across four domains, we curate our dataset Logic-Induced-Long-Tail(LINT) that consists of 54K long-tail knowledge statements. We also release 54K head distribution statements that are also searched with the LINK framework. Domain-wise statistics of LINT in Appendix H.
3 Generating Long-tail Inferential Knowledge with LINK
In this section, we compare LINK’s ability to generate long-tail inferential knowledge with instruction-only LLMs, ChatGPT and GPT4, who do not use knowledge beam search.
3.1 Instruction-Only Knowledge Generation
Symbolic Rule    is_allergic_to(Person X, Food allergen A)& is_ ingredient_in(Ingredient Z, Name of a dish or food B) & is_one_type_of(Ingredient Z, Food allergen A) $\to$ is_not_able_to_eat(Person P, Name of a dish or food B)
Prompt In the following sentence, A is a Food allergen, B is a Name of a dish or food, Z is a Ingredient. Find values of A, B, Z to fill in the blank in the sentence ‘If Person X is allergic to [A] and [ Z] is a ingredient in [ B] and [Z] is one type of [ A], then Person X is not able to eat [ B].’ and make it a grammatical and correct sentence. Give me 50 values in the format ‘1. A=, B=, Z=’.